# HC-Mamba: Vision MAMBA with Hybrid Convolutional Techniques for Medical Image Segmentation

### v2

@@ 1 Introduction
We conducted extensive performance evaluations of the proposed model. The results show that our model has high accuracy (94.84%), mIoU (80.60%) and validity of DSC (89.25%).
@@ 3.2 Results
For the ISIC2017 and ISIC2018 datasets, HC-Mamba performs well on mIoU and Dice compared to other models. Specifically, HC-Mamba has a 1.46% and 1% advantage over MedMamba on mIoU and Dice, respectively, while it has a 2.74% and 1.7% advantage over Unet on mIoU and Dice.
@@
| HC-Mamba | 80.60 | 89.25 | 94.84 | 97.08 | 87.90 |
For the Synapse datasets, HC-Mamba performs well on DSC and HD95 compared to other models. Specifically, HC-Mamba has a 0.89% and 2.48% advantage over MedMamba and VM-Unet on Dice.

### v3

@@ 1 Introduction
We conducted extensive performance evaluations of the proposed model. The results show that our model has high accuracy (94.84%), mIoU (80.72%) and validity of DSC (89.26%).
@@ 2.2.2 HC-SSM Block
2.2.3 Loss function
In the field of medical image segmentation, the most crucial evaluation metrics are the overlap between the segmentation results and the ground truth, the accuracy of the boundaries, and the similarity. Therefore, we designed a comprehensive weighted loss function that combines mIoU loss, Dice loss, and Boundary loss. The rationale for selecting these three loss functions is that each optimizes a distinct aspect of the segmentation task. The mIoU loss function evaluates the overlap of the segmented area, the Dice loss function measures the similarity of the segmented area, and the boundary loss function focuses on the accuracy of the segmentation boundaries. The combination of these three losses allows for a comprehensive optimization of the performance of the segmentation model.
Specifically, the loss function is expressed as follows:
where , , and represent the mIoU loss, Dice loss, and boundary loss, respectively. The terms , , and are the corresponding weight coefficients. The term and denote the boundary point sets of the predicted segmentation and the ground truth segmentation, respectively, and represents the Euclidean distance between point and point .
The comprehensive loss function, which is the result of weighting and combining the three loss functions, is able to simultaneously optimize the overlap, similarity, and boundary accuracy of the segmentation regions,thereby improving the overall segmentation performance.
@@ 3.2 Results
For the ISIC2017 and ISIC2018 datasets, HC-Mamba performs well on mIoU and Dice compared to other models. Specifically, HC-Mamba has a 1.59% and 0.9% advantage over MedMamba on mIoU and Dice, respectively, while it has a 2.84% and 1.71% advantage over Unet on mIoU and Dice.
@@
| HC-Mamba | 80.72 | 89.26 | 94.84 | 97.08 | 88.90 |
For the Synapse datasets, HC-Mamba performs well on DSC and HD95 compared to other models. Specifically, HC-Mamba has a 0.91% and 2.5% advantage over MedMamba and VM-Unet on Dice.